# Adaptive sampling accelerates the hybrid deviational particle simulations

### v1

@@ 4 Numerical experiments
This section conducts several typical numerical experiments on the 1-D spatial and 3-D velocity coordinates and makes a detailed comparison between HDP-AS and HDP. We set the spatial domain to be $x\in\left[0,4\pi\right]$ with periodic boundary condition and choose $N_{x}=100$ grids with the spatial spacing $\Delta x=\frac{4\pi}{N_{x}}$. The time step is set as $\Delta t=\Delta x/10$. The partitioning parameters for the adaptive reconstruction in Algorithm 4 are set as $\vartheta=0.1$, $q=16$ unless otherwise specified. We control the numerical particle numbers by the parameters $N_{\text{eff }}$ and $N_{\text{eff }}^{F}$, and let
$N_{f,k}=\frac{\rho(x_{k},t)}{N_{\text{eff }}^{F}},\quad N_{p,k}=\frac{\rho_{p}(x_{k},t)}{N_{\text{eff }}},\quad N_{n,k}=\frac{\rho_{n}(x_{k},t)}{N_{\text{eff }}},\quad k=1,\dots,N_{x}.$ (27)
In order to visualize the high-dimensional numerical solutions, we adopt the following 1-D and 2-D projections
@@ 4.1 Linear Landau damping
where the constant $\alpha$ characterizes the amplitude of the perturbation. We conduct the tests until the final time $t_{fin}=5$. Table 1 presents the efficiency tests for $\alpha=0.001,0.01,0.1$ and different sample sizes. The HDP-AS method runs roughly ten times faster than the HDP method, along with smaller errors. For example, when $\alpha=0.01$, $N_{\text{eff }}=5\times 10^{-6}$ and $N_{\text{eff }}^{F}=1\times 10^{-5}$, HDP-AS takes 403.26 seconds while HDP takes 4626.91 seconds, achieving an acceleration of 11.47 times. In terms of the accuracy, the relative $L^{2}$ error $\mathcal{E}_{2}[p_{2d}](t_{fin})$ of HDP-AS (0.0264) is also preferable to that of HDP (0.0451). Figure 3 displays the snapshots of $p_{2d}(x,v_{1},t_{fin})$ produced by HDP and HDP-AS corresponding to this group of parameters and as excepted, a good agreement can be readily observed. Figure 4 displays the histogram of the acceptance rates in HDP. More than 73% of the acceptance rates are less than 0.01. The low acceptance rate leads to a high sampling cost. On the contrary, HDP-AS achieves high-dimensional sampling without rejection by approximating the source term $\Delta m_{k}(\bm{v})$ with an adaptive piecewise constant function $\overline{\Delta m_{k}}(\bm{v})$ defined in Eq. (18). Figure 5 plots $\int_{\mathbb{R}^{3}}\overline{\Delta m_{k}}(\bm{v})\textup{d}v_{3}$, the projection of adaptive piecewise constant approximation $\overline{\Delta m_{k}}(\bm{v})$ in $v_{1}v_{2}$-plane with $k=50$. In accordance with the changes of the distribution $f(x,\bm{v},t)$ (see Figure 3 for the 2-D projections in $v_{1}x$-plane), this projection exhibits significant changes in the region $(v_{1},v_{2})\in\left[-2,2\right]\times\left[-2,2\right]$, which reflects the intrinsic adaptivity of HDP-AS. Just because of the adaptivity of $\overline{\Delta m_{k}}(\bm{v})$, new deviational particles sampled by HDP-AS are of higher quality than those by HDP and thus able to capture the characteristics of $\Delta m_{k}(\bm{v})$ accurately with a smaller size. For instance, HDP-AS samples $177$ new P-particles and $166$ new N-particles, whereas HDP samples $412$ new P-particles and $370$ new N-particles for $\alpha=0.01$, $N_{\text{eff }}=5\times 10^{-6}$, $N_{\text{eff }}^{F}=1\times 10^{-5}$ and $t=1$ in this example. Much less new deviational particles being added may enable other remaining steps such as simulating collisions via TA and transport to take less time (see Figure 1).
@@ 5 Conclusion and discussion
Owing to the above two improvements, the proposed HDP-AS method successfully reduces the computational cost of the HDP method. There remain some aspects to be explored, such as the error analysis of HDP-AS, the application to practical problems in higher dimensions, and whether more efficient uniformity measure manner can further accelerate the reconstruction of the adaptive piecewise constant approximation.

### v2

@@ 4 Numerical experiments
This section conducts five typical numerical experiments on the 1-D spatial and 3-D velocity coordinates, including the linear and nonlinear Landau damping problems, two stream instability problem, Bump on Tail problem and Rosenbluth’s problem. A careful benchmark of HDP on the linear and nonlinear Landau damping against PIC-DSMC [3], which combines a Particle in Cell (PIC) method on the advection and a Direct Simulation Monte Carlo (DSMC) method on the collisions, as well as on Bump on Tail problem and Rosenbluth’s problem against the TA method [12], has been conducted and demonstrates the accuracy and efficiency of HDP. Therefore this section only focuses on a clear and one-to-one comparison between HDP-AS and HDP. We set the spatial domain to be $x\in\left[0,4\pi\right]$ with periodic boundary condition and choose $N_{x}=100$ grids with the spatial spacing $\Delta x=\frac{4\pi}{N_{x}}$. The time step is set as $\Delta t=\Delta x/10$. The partitioning parameters for the adaptive reconstruction in Algorithm 4 are set as $\vartheta=0.1$, $q=16$ unless otherwise specified. We control the numerical particle numbers by the parameters $N_{\text{eff }}$ and $N_{\text{eff }}^{F}$. In order to visualize the high-dimensional numerical solutions, we adopt the following 1-D and 2-D projections
@@ 4.1 Linear Landau damping
where the constant $\alpha$ characterizes the amplitude of the perturbation. We conduct the tests until the final time $t_{fin}=5$. Table 1 presents the efficiency tests for $\alpha=0.001,0.01,0.1$ and different sample sizes. The HDP-AS method runs roughly ten times faster than the HDP method, along with smaller errors. For example, when $\alpha=0.01$, $N_{\text{eff }}=5\times 10^{-6}$ and $N_{\text{eff }}^{F}=1\times 10^{-5}$, HDP-AS takes 403.26 seconds while HDP takes 4626.91 seconds, achieving an acceleration of 11.47 times. In terms of the accuracy, the relative $L^{2}$ error $\mathcal{E}_{2}[p_{2d}](t_{fin})$ of HDP-AS (0.0264) is also preferable to that of HDP (0.0451). Figure 3 displays the snapshots of $p_{2d}(x,v_{1},t_{fin})$ produced by HDP and HDP-AS corresponding to this group of parameters and as excepted, a good agreement can be readily observed. Figure 4 displays the histogram of the acceptance rates in HDP. More than 73% of the acceptance rates are less than 0.01. The low acceptance rate leads to a high sampling cost. On the contrary, HDP-AS achieves high-dimensional sampling without rejection by approximating the source term $\Delta m_{k}(\bm{v})$ with an adaptive piecewise constant function $\overline{\Delta m_{k}}(\bm{v})$ defined in Eq. (19). Figure 5 plots $\int_{\mathbb{R}^{3}}\overline{\Delta m_{k}}(\bm{v})\textup{d}v_{3}$, the projection of adaptive piecewise constant approximation $\overline{\Delta m_{k}}(\bm{v})$ in $v_{1}v_{2}$-plane with $k=50$. In accordance with the changes of the distribution $f(x,\bm{v},t)$ (see Figure 3 for the 2-D projections in $v_{1}x$-plane), this projection exhibits significant changes in the region $(v_{1},v_{2})\in\left[-2,2\right]\times\left[-2,2\right]$, which reflects the intrinsic adaptivity of HDP-AS. Just because of the adaptivity of $\overline{\Delta m_{k}}(\bm{v})$, new deviational particles sampled by HDP-AS are of higher quality than those by HDP and thus able to capture the characteristics of $\Delta m_{k}(\bm{v})$ accurately with a smaller size. For instance, HDP-AS samples $177$ new P-particles and $166$ new N-particles, whereas HDP samples $412$ new P-particles and $370$ new N-particles for $\alpha=0.01$, $N_{\text{eff }}=5\times 10^{-6}$, $N_{\text{eff }}^{F}=1\times 10^{-5}$ and $t=1$ in this example. Much less new deviational particles being added may enable other remaining steps such as simulating collisions via TA and advection to take less time (see Figure 1).
@@ 5 Conclusion and discussion
Owing to the above two improvements, the proposed HDP-AS method successfully reduces the computational cost of the HDP method. The same idea could be applied to other sampling tasks, provided that each part of the function to be sampled has already been represented by multiple particle sets (e.g., in this work, the $Q(f_{p},m)$ and $Q(f_{n},m)$ in $\Delta m_{k}(\bm{v})_{\pm}$ are represented by particle sets $\mathbb{M}^{+}_{k}$ and $\mathbb{M}^{-}_{k}$ in Eq. (15)). There remain some aspects to be explored, such as the error analysis of HDP-AS, the application to practical problems in higher dimensions, and whether more efficient uniformity measure manner can further accelerate the reconstruction of the adaptive piecewise constant approximation.